OpenStudy (anonymous):

help:) solving by competing the square: 3x^2-10x+8=0

OpenStudy (anonymous):

Divide both sides by 3: x^2-(10 x)/3+8/3 = 0 Subtract 8/3 from both sides: x^2-(10 x)/3 = -8/3 Add 25/9 to both sides: x^2-(10 x)/3+25/9 = 1/9 Factor the left hand side: (x-5/3)^2 = 1/9 Take the square root of both sides: sqroot(x-5/3) = 1/3 Eliminate the absolute value: x-5/3 = -1/3 or x-5/3 = 1/3 Add 5/3 to both sides: x = 4/3 or x-5/3 = 1/3 Add 5/3 to both sides: x = 4/3 or x = 2

OpenStudy (anonymous):

\[3(x^2-\frac{10}{3}x)=-8\] \[(\frac{10}{3})^2=\frac{100}{9}......3(x^2-\frac{10}{3}x + \frac{100}{9} )=-8+\frac{300}{9}\] \[3(x-\frac{10}{3})^2=\frac{76}{3}............\sqrt{(x-\frac{10}{3})}=\frac{\sqrt{76}}{3}\]\[x=\frac{10+/-2\sqrt{19}}{3}\]